OpenStudy (anonymous):

How do you get the relative maximum and minimum values of a x3 graph that does not have definite curves? For instance, in the graph of x^3-6x^2=15 , the graph looks like a traditional x^3 graph

OpenStudy (anonymous):

first derivative = 0

OpenStudy (anonymous):

Is it even possible to have a relative maximum or minimum for this type of graph? When you plug it into the clculator you can see that the curves are connected, which takes away the possibility of getting an accurate point

OpenStudy (anonymous):

if y=x^3-6x^2-15 y'=3x^2-12x which will gives you 2 critical points: x=0 & x=4 I think (without checking) x=0 is local max check x=4

OpenStudy (anonymous):

x=4 should be local min

OpenStudy (anonymous):

here is the graph - you can see that x=0 is local max & x=4 is loc min

OpenStudy (anonymous):

sorry I had typo, it's + 15

OpenStudy (anonymous):

This is what the graph looks like

OpenStudy (anonymous):

how do you find it with that?

OpenStudy (anonymous):

this one : y=x^3 - 6x^2 +15 see attached - still has local max @ x=0 & local min @ x=4

OpenStudy (anonymous):

I took first derivative: y=3x^2 -12x critical points when y'=0 solve it for x

OpenStudy (mathteacher1729):

Go to http://www.wolframalpha.com and then type extrema of x^3-6x^2

OpenStudy (anonymous):

that one looks like y=x^3 y'=3x^2 y'=0 when x=0 it's not min or max, it's called "Inflection point" see: http://en.wikipedia.org/wiki/Inflection_point

OpenStudy (anonymous):

good?

OpenStudy (anonymous):

awesome